how do i y=2x and x+y=12 by solving system of equations by the substitution method?
Answer by stanbon(75887)

y=2x and x+y=12
----
Substitute for "y" and solve for "x":
x+(2x) = 12
3x = 12
x = 4
----
Solve for "y":
y = 2x
y = 2*4
y = 8
====================
Cheers,
Stan H.
====================